# Dark Matter (DM); Dark Stars (DSs); High-z galaxies; Supermassive Black Holes (SMBHs), Little Red Dots (LRDs); Blue Monsters; Over-massive BH galaxies (OBGs)

### v1

@@ Appendix B Dark Stars
In this Appendix we expand upon the material presented Sec. 2 of the main body of the paper. The effects of Dark Matter (DM) heating on the formation of the first stars in the universe were first analyzed in Ref. Spolyar:2008dark, who found the conditions necessary for the formation of stars powered by annihilations of Weakly Interacting Massive Particles (WIMPs). [^8] Note that the same WIMP models that could explain the AMS-02 and Fermi anomalies could also power Dark Stars. Namely: i) energy for Dark Matter annihilations can be deposited efficiently in the protostellar gas clouds, and ii) Dark Matter heating overcomes all available cooling mechanisms for the collapsing baryonic cloud. For zero metallicity clouds, such as those out of which Dark Stars could form, the cooling is relatively inefficient, and is provided by $H_{2}$ cooling, supplemented by $H$ line and Compton cooling. The rate, per unit volume, at which energy from DM annihilations is deposited is:
@@
As protostellar molecular clouds collapse, the gravitational potential steepens, and, as such, DM orbits will respond, leading to an enhancement of the DM densities at the center of DM halos where the first stars form. The Blumenthal (Blumenthal:1985) method of Adiabatic Contraction (AC) is used in Ref Spolyar:2008dark, who find that the DM density, at the edge of the baryonic core, tracks the baryonic number density ($n_{B}$) in the following approximate manner:
@@
wheres outside the core $\rho_{DM}\propto r^{-1.9}$. Those results were subsequently confirmed (to within factors of a few) by more precise calculations in which the possibility of non circular DM orbits is included, and various initial DM profiles, in addition to the standard NFW, are explored (Freese:2008dmdens). When compared to an initial DM profile, the DM densities can be enhanced by as much as four to five orders of magnitude via the process described above. This leads to an enhancement of 8-10 orders of magnitude of the heating rate (see Eq. 2). At the same time, as the baryonic density becomes larger, an appreciable fraction of the energy released in DM annihilation products can be deposited inside the cloud (Spolyar:2008dark).
@@
Ref. Spolyar:2008dark finds that for $100$ GeV WIMPs the DM heating rate balances all cooling mechanisms whenever the baryonic number density $n_{B}\sim 10^{13}\mathrm{cm}^{-3}$. For lighter/heavier WIMPs this happens at at lower/higher values of $n_{B}$, in view of their more/less efficient DM heating rates (see Eq. 2). Soon after DM heating overtakes cooling, the collapse will halt, and a proto Dark Star is formed. For instance, for $100$ GeV WIMPs, this happens when $n_{B}\sim 10^{17}\mathrm{cm}^{-3}$, the size of the baryonic core is of the order of $\sim 17$ A.U.s and it has a mass of $\sim 0.6M_{\odot}$. At that stage the luminosity due to DM heating is $L_{DM}\sim 140L_{\odot}$. Therefore, the standard picture of the formation of the first stars is drastically modified by WIMP Dark Matter annihilations which could halt the baryonic cloud collapse well before it reaches the critical temperatures to fuse Hydrogen.
@@
Figure 8: HR diagram for dark stars for accretion rate $\dot{M}=10^{-3}M_{\odot}$/yr. and a variety of WIMP masses as labeled for the two cases: (i) “without capture” but with extended adiabatic contraction (dotted lines) and (ii) “with capture” (solid lines). The case with capture is for product of scattering cross section times ambient WIMP density $\sigma_{c}\bar{\rho}_{\chi}=10^{-39}{\rm cm}^{2}\times 10^{13}$GeV/cm3. Also labeled are stellar masses reached by the DS on its way to becoming supermassive. The final DS mass was taken to be $1.5\times 10^{5}M_{\odot}$, but SDMSs could grow in principle past this mass. (Figure adapted from (Freese:2010smds).)
@@
In Fig. 8 we reproduce the HR diagram for Dark Stars powered by WIMPs of three different masses growing at an accretion rate of $10^{-3}M_{\odot}\mathrm{yr}^{-1}$ from Freese:2010smds, who model using polytropes of variable index the evolution of Dark Stars from their formation ($\sim 1M_{\odot}$) until they become supermassive ($M\gtrsim 10^{5}M_{\odot}$). Note how Dark Stars formed via the extended AC mechanism (labeled in the figure Without Capture) are cooler than those that formed via the DM capture mechanism (labeled With Capture). This is because the later underwent a Kelviln-Helmholtz contraction phase (see horizontal solid lines in the HR diagram) once they ran out of the initial AC DM reservoir. Note how lighter WIMPs are more efficient in heating DSs. This is to be expected from Eq. 2, and can be explicitly seen in the HR diagram, as at the same stellar mass, a DS powered by a lighter WIMP is puffier.
@@
WIMPs are not the only DM particle models that can power Dark Stars. For instance, Ref. Wu:2022SIDMDS identified a specific model of Self Interacting Dark Matter (SIDM) which could satisfy bounds from CMB, direct and indirect detection, and, at the same time, deposit energy from annihilations into protostellar zero metallicity gas clouds efficiently. Ongoing preliminary work on light Dark Matter models that anihilate via number changin processes (e.g. SIMPs and Co-SIMPs) indicates that DSs could be powered those kind of DM particles as well. Therefore, DS are not restricted to the WIMP landscape, and we plan to explore the full DM parameter space that allows for DS to form.
@@ Appendix C Dark Stars and DCBHs: Oversized BH Seeds
Since its initial theoretical proposal by Belgman:2006, the DCBH scenario has undergone several refinements and, in view of recent data has sparked significant interest. In a nutshell, cool, rapidly accreting, supermassive Population III stars can seed Direct Collapse Black Holes. Finding environments where such rapid accretion rates are possible is challenging, but not impossible, according to simulations. For instance, in atomic cooling halos, baryonic infall rates range in the relatively wide $10^{-2}-10^{1}M_{\odot}/yr$ range (Latif:2013pyq; Choon:2018; Patrick_2023; Toyouchi_2022), although typically they stay less than $1M_{\odot}\mathrm{yr}^{-1}$. For accretion rates exceeding $\sim 0.1M_{\odot}\mathrm{yr}^{-1}$, Pop III stars enter a supergiant phase, leading to negligible production of ionizing radiation (e.g. Hosokawa_2012). Thus accretion can proceed without much feedback. Once they grow supermassive, Pop III stars can even accrete at rates exceeding the atomic cooling limit ($10M_{\odot}\mathrm{yr}^{-1}$), as shown by Ref. Haemmerle:2019. DCBHs are believed to form inside atomically cooled DM halos (i.e. $M_{halo}\gtrsim 10^{7}M_{\odot}$) and only if there is a mechanism to destroy molecular hydrogen, or suppress its formation. Lyman-Werner (LW) photons from a nearby starburst galaxy (e.g. Belgman:2006) or collisional dissociation (e.g. Kiyuna_2023) are two of the most common DCBH formation channels. In addition, dynamically heated atomically cooled halos could be sites for DCBHs even if $H_{2}$ is not fully dissociated (Wise_2019; Li__2021; Toyouchi_2022). In view of the very special conditions required to be met, the fraction of high redshift DM halos expected to contain a DCBH is quite rare (Natarajan:2017).
@@
Dark Stars are not hot enough to emit strong ionizing fluxes that would completely shut off accretion. As such, they could potentially accrete the entirety of the gas in the surrounding cloud/disk (e.g. Freese:2010smds; Banik:2019). The Jeans scale for $H_{2}$ cooled gas clouds is $\sim 1000M_{\odot}$, and it would seem that this is a reasonable guess for the upper mass of Dark Stars formed in microhalos where $H_{2}$ cooling is dominant. However, Dark Matter heating itself leads to a suppression of the fragmentation of the gas cloud (Stacy:2014), so, even in the absence of any mechanism that would increase the supply of gravitationally unstable gas  [^9] Such a mechanism is required in the DCBH scenario. Dark Stars can easily grow to masses exceeding $1000M_{\odot}$ (Freese:2010smds), thus providing natural seeds for massive or even supermassive BHs. Additionally, the formation of SMDSs does not imply the need for any mechanism to destroy $H_{2}$. In fact Dark Stars can form both in atomically and $H_{2}$ cooled halos (Spolyar:2009). Moreover, when contrasted against DCBHs, SMDSs are able to form over a wider variety of redshifts, and a sizable fraction of DM halos would be able to host SMDS. In contrast, only a small fraction of high-z halos would be able to host DCBHs. Lastly, unlike the supermassive Population III stellar precursors of DCBHs, Supermassive Dark Stars do not require very high accretion rates to remain cool. Therefore, in view of all the points above, an SMDS seed is a more flexible model for the formation of the most massive high-z quasars.
@@ Appendix D Collapse of SMDSs via mergers of halos
First, any individual star moving in a field of stars experiences dynamical friction (Chandrasekhar:1943DynFric). In this collisionless medium case, the star is decelerated due to the fluctuating gravitational field in which it moves. In our merger scenario, this would be relevant if the halo with which the DS host halo merges will have formed stars before the mergers (see Fig. 7). Note that the formalism developed in (Chandrasekhar:1943DynFric) also covers the case of dynamical friction when a star is moving with respect to a Dark Matter halo, as would be the case for a SMDS host halo merging with another DM halo, irrespective of its contents. Second, a massive star moving in a gaseous medium can also experience dynamical friction, as a consequence of the gravitational attraction between the massive star and its wake in the ambient medium (Ostriker:1999DynFricGas). In our merger scenario, if the DS hosting halo merges with a cold DM halo (in which the gas clouds have not yet collapsed to the point of forming stars), this type of dynamical friction could be relevant. Lastly, the companion halo could host a very massive central object itself, such as a Black Hole that has grown via accretion (or even another SMDS). During these types of mergers, massive central objects usually orbit each other, since they lose sufficient energy to dynamical friction [^10] In this scenario dynamical friction can be caused by either stars, gas, or colisionless Dark Matter. and thus become gravitationally bound (e.g. Chen:2022DynFrictionMergers; Tremmel:2015DynFricMergers). Typically they start orbiting when separated by as much as a few tens of kpc, lose energy (i.e. orbits shrinking) via dynamical friction, and eventually merge (e.g. Tremmel:2015DynFricMergers). Thus, during those orbits each of the two massive objects would no longer be perfectly aligned with the center of its respective DM halo, or with the center of the newly formed merged DM halo.
@@
We estimate below the effects of moving a SMDSs about $10A.U.$s from the its location at center of its host DM halo. This corresponds to shifting the SMDSs by a distance comparable to its radius. DM heating is controlled by the square of the DM density (see Eq. 2). To a good approximation the DM density in the vicinity of the SMDSs scales with distance from the DM halo center as $r^{-1.9}$ (Spolyar:2008dark; Freese:2008dmdens). Therefore, the DM heating inside of the Dark Star is reduced by a factor of $(20/10)^{-3.8}\simeq 0.07$. In other words, the DM now provides less than $10\%$ of the heating needed to sustain the SMDSs. Of course, during the two halo merger, the DM densities can be enhanced as well. Assuming those two halos are similar, this enhancement would be by a factor of $\sim 2$. Thus, in this case the DM heating is reduced by a factor of $2^{2}\times 0.07\simeq 0.28$, which is still quite important.
@@
This significant reduction in the DM heating should lead to a collapse of the SMDSs to a SMBH, since such massive objects, once the fuel is reduced, collapse directly to Black Holes. As a consequence of the merger, in addition to the collapse of the SMDS to a SMBH, an episode of rapid star formation, i.e. a starburst, is triggered in the vicinity of the SMBH (e.g. Silk:2005; Levin:2007MNRAS; carilli:2013; heckman:2014; MF2023). Moreover, the reservoir of baryons in the vicinity of the SMDS is replenished now via the merger. Thus, the SMBH can continue to grow at the Eddington rate, and be observed as the X-ray portion of the UHZ1 radiation. The stellar population embedded in those two merging halos is responsible for the majority of the IR observed data. In Fig. 7 we present a schematic representation of the scenario discussed above. Or, there could already be one halo in which the SMDS has already collapsed to a SMBH, and that halo then merges with another. The merger then creates a starburst in addition to the SMBH that was already there.

### v2

@@ Appendix B Dark Stars
In this Appendix we expand upon the material presented Sec. 2 of the main body of the paper. The effects of Dark Matter (DM) heating on the formation of the first stars in the universe were first analyzed in Ref. Spolyar et al. (2008), who found the conditions necessary for the formation of stars powered by annihilations of Weakly Interacting Massive Particles (WIMPs). [^9] Note that the same WIMP models that could explain the AMS-02 and Fermi anomalies could also power Dark Stars. Namely: i) energy for Dark Matter annihilations can be deposited efficiently in the protostellar gas clouds, and ii) Dark Matter heating overcomes all available cooling mechanisms for the collapsing baryonic cloud. For zero metallicity clouds, such as those out of which Dark Stars could form, the cooling is relatively inefficient, and is provided by $H_{2}$ cooling, supplemented by $H$ line and Compton cooling. The rate, per unit volume, at which energy from DM annihilations is deposited is:
@@
As protostellar molecular clouds collapse, the gravitational potential steepens, and, as such, DM orbits will respond, leading to an enhancement of the DM densities at the center of DM halos where the first stars form. The Blumenthal (Blumenthal et al., 1986) method of Adiabatic Contraction (AC) is used in Ref Spolyar et al. (2008), who find that the DM density, at the edge of the baryonic core, tracks the baryonic number density ($n_{B}$) in the following approximate manner:
@@
wheres outside the core $\rho_{DM}\propto r^{-1.9}$. Those results were subsequently confirmed (to within factors of a few) by more precise calculations in which the possibility of non circular DM orbits is included, and various initial DM profiles, in addition to the standard NFW, are explored (Freese et al., 2009). When compared to an initial DM profile, the DM densities can be enhanced by as much as four to five orders of magnitude via the process described above. This leads to an enhancement of 8-10 orders of magnitude of the heating rate (see Eq. 2). At the same time, as the baryonic density becomes larger, an appreciable fraction of the energy released in DM annihilation products can be deposited inside the cloud (Spolyar et al., 2008).
@@
Ref. Spolyar et al. (2008) finds that for $100$ GeV WIMPs the DM heating rate balances all cooling mechanisms whenever the baryonic number density $n_{B}\sim 10^{13}\mathrm{cm}^{-3}$. For lighter/heavier WIMPs this happens at at lower/higher values of $n_{B}$, in view of their more/less efficient DM heating rates (see Eq. 2). Soon after DM heating overtakes cooling, the collapse will halt, and a proto Dark Star is formed. For instance, for $100$ GeV WIMPs, this happens when $n_{B}\sim 10^{17}\mathrm{cm}^{-3}$, the size of the baryonic core is of the order of $\sim 17$ A.U.s and it has a mass of $\sim 0.6M_{\odot}$. At that stage the luminosity due to DM heating is $L_{DM}\sim 140L_{\odot}$. Therefore, the standard picture of the formation of the first stars is drastically modified by WIMP Dark Matter annihilations which could halt the baryonic cloud collapse well before it reaches the critical temperatures to fuse Hydrogen.
@@
Figure 8: HR diagram for dark stars for accretion rate $\dot{M}=10^{-3}M_{\odot}$/yr. and a variety of WIMP masses as labeled for the two cases: (i) “without capture” but with extended adiabatic contraction (dotted lines) and (ii) “with capture” (solid lines). The case with capture is for product of scattering cross section times ambient WIMP density $\sigma_{c}\bar{\rho}_{\chi}=10^{-39}{\rm cm}^{2}\times 10^{13}$GeV/cm3. Also labeled are stellar masses reached by the DS on its way to becoming supermassive. The final DS mass was taken to be $1.5\times 10^{5}M_{\odot}$, but SDMSs could grow in principle past this mass. (Figure adapted from (Freese et al., 2010).)
@@
In Fig. 8 we reproduce the HR diagram for Dark Stars powered by WIMPs of three different masses growing at an accretion rate of $10^{-3}M_{\odot}\mathrm{yr}^{-1}$ from Freese et al. (2010), who model using polytropes of variable index the evolution of Dark Stars from their formation ($\sim 1M_{\odot}$) until they become supermassive ($M\gtrsim 10^{5}M_{\odot}$). Note how Dark Stars formed via the extended AC mechanism (labeled in the figure Without Capture) are cooler than those that formed via the DM capture mechanism (labeled With Capture). This is because the later underwent a Kelviln-Helmholtz contraction phase (see horizontal solid lines in the HR diagram) once they ran out of the initial AC DM reservoir. Note how lighter WIMPs are more efficient in heating DSs. This is to be expected from Eq. 2, and can be explicitly seen in the HR diagram, as at the same stellar mass, a DS powered by a lighter WIMP is puffier.
@@
WIMPs are not the only DM particle models that can power Dark Stars. For instance, Ref. Wu et al. (2022) identified a specific model of Self Interacting Dark Matter (SIDM) which could satisfy bounds from CMB, direct and indirect detection, and, at the same time, deposit energy from annihilations into protostellar zero metallicity gas clouds efficiently. Ongoing preliminary work on light Dark Matter models that anihilate via number changin processes (e.g. SIMPs and Co-SIMPs) indicates that DSs could be powered those kind of DM particles as well. Therefore, DS are not restricted to the WIMP landscape, and we plan to explore the full DM parameter space that allows for DS to form.
@@ Appendix C Dark Stars and DCBHs: Oversized BH Seeds
Since its initial theoretical proposal by Begelman et al. (2006), the DCBH scenario has undergone several refinements and, in view of recent data has sparked significant interest. In a nutshell, cool, rapidly accreting, supermassive Population III stars can seed Direct Collapse Black Holes. Finding environments where such rapid accretion rates are possible is challenging, but not impossible, according to simulations. For instance, in atomic cooling halos, baryonic infall rates range in the relatively wide $10^{-2}-10^{1}M_{\odot}/yr$ range (Latif et al., 2013; Chon et al., 2018; Patrick et al., 2023; Toyouchi et al., 2022), although typically they stay less than $1M_{\odot}\mathrm{yr}^{-1}$. For accretion rates exceeding $\sim 0.1M_{\odot}\mathrm{yr}^{-1}$, Pop III stars enter a supergiant phase, leading to negligible production of ionizing radiation (e.g. Hosokawa et al., 2012). Thus accretion can proceed without much feedback. Once they grow supermassive, Pop III stars can even accrete at rates exceeding the atomic cooling limit ($10M_{\odot}\mathrm{yr}^{-1}$), as shown by Ref. Haemmerlé et al. (2019). DCBHs are believed to form inside atomically cooled DM halos (i.e. $M_{halo}\gtrsim 10^{7}M_{\odot}$) and only if there is a mechanism to destroy molecular hydrogen, or suppress its formation. Lyman-Werner (LW) photons from a nearby starburst galaxy (e.g. Begelman et al., 2006) or collisional dissociation (e.g. Kiyuna et al., 2023) are two of the most common DCBH formation channels. In addition, dynamically heated atomically cooled halos could be sites for DCBHs even if $H_{2}$ is not fully dissociated (Wise et al., 2019; Li et al., 2021; Toyouchi et al., 2022). In view of the very special conditions required to be met, the fraction of high redshift DM halos expected to contain a DCBH is quite rare (Natarajan et al., 2017).
@@
Dark Stars are not hot enough to emit strong ionizing fluxes that would completely shut off accretion. As such, they could potentially accrete the entirety of the gas in the surrounding cloud/disk (e.g. Freese et al., 2010; Banik et al., 2019). The Jeans scale for $H_{2}$ cooled gas clouds is $\sim 1000M_{\odot}$, and it would seem that this is a reasonable guess for the upper mass of Dark Stars formed in microhalos where $H_{2}$ cooling is dominant. However, Dark Matter heating itself leads to a suppression of the fragmentation of the gas cloud (Stacy et al., 2014), so, even in the absence of any mechanism that would increase the supply of gravitationally unstable gas  [^10] Such a mechanism is required in the DCBH scenario. Dark Stars can easily grow to masses exceeding $1000M_{\odot}$ (Freese et al., 2010), thus providing natural seeds for massive or even supermassive BHs. Additionally, the formation of SMDSs does not imply the need for any mechanism to destroy $H_{2}$. In fact Dark Stars can form both in atomically and $H_{2}$ cooled halos (Spolyar et al., 2009). Moreover, when contrasted against DCBHs, SMDSs are able to form over a wider variety of redshifts, and a sizable fraction of DM halos would be able to host SMDS. In contrast, only a small fraction of high-z halos would be able to host DCBHs. Lastly, unlike the supermassive Population III stellar precursors of DCBHs, Supermassive Dark Stars do not require very high accretion rates to remain cool. Therefore, in view of all the points above, an SMDS seed is a more flexible model for the formation of the most massive high-z quasars.
@@ Appendix D Collapse of SMDSs via mergers of halos
First, any individual star moving in a field of stars experiences dynamical friction (Chandrasekhar, 1943). In this collisionless medium case, the star is decelerated due to the fluctuating gravitational field in which it moves. In our merger scenario, this would be relevant if the halo with which the DS host halo merges will have formed stars before the mergers (see Fig. 7). Note that the formalism developed in (Chandrasekhar, 1943) also covers the case of dynamical friction when a star is moving with respect to a Dark Matter halo, as would be the case for a SMDS host halo merging with another DM halo, irrespective of its contents. Second, a massive star moving in a gaseous medium can also experience dynamical friction, as a consequence of the gravitational attraction between the massive star and its wake in the ambient medium (Ostriker, 1999). In our merger scenario, if the DS hosting halo merges with a cold DM halo (in which the gas clouds have not yet collapsed to the point of forming stars), this type of dynamical friction could be relevant. Lastly, the companion halo could host a very massive central object itself, such as a Black Hole that has grown via accretion (or even another SMDS). During these types of mergers, massive central objects usually orbit each other, since they lose sufficient energy to dynamical friction [^11] In this scenario dynamical friction can be caused by either stars, gas, or colisionless Dark Matter. and thus become gravitationally bound (e.g. Chen et al., 2022; Tremmel et al., 2015). Typically they start orbiting when separated by as much as a few tens of kpc, lose energy (i.e. orbits shrinking) via dynamical friction, and eventually merge (e.g. Tremmel et al., 2015). Thus, during those orbits each of the two massive objects would no longer be perfectly aligned with the center of its respective DM halo, or with the center of the newly formed merged DM halo.
@@
We estimate below the effects of moving a SMDSs about $10A.U.$s from the its location at center of its host DM halo. This corresponds to shifting the SMDSs by a distance comparable to its radius. DM heating is controlled by the square of the DM density (see Eq. 2). To a good approximation the DM density in the vicinity of the SMDSs scales with distance from the DM halo center as $r^{-1.9}$ (Spolyar et al., 2008; Freese et al., 2009). Therefore, the DM heating inside of the Dark Star is reduced by a factor of $(20/10)^{-3.8}\simeq 0.07$. In other words, the DM now provides less than $10\%$ of the heating needed to sustain the SMDSs. Of course, during the two halo merger, the DM densities can be enhanced as well. Assuming those two halos are similar, this enhancement would be by a factor of $\sim 2$. Thus, in this case the DM heating is reduced by a factor of $2^{2}\times 0.07\simeq 0.28$, which is still quite important.
@@
This significant reduction in the DM heating should lead to a collapse of the SMDSs to a SMBH, since such massive objects, once the fuel is reduced, collapse directly to Black Holes. As a consequence of the merger, in addition to the collapse of the SMDS to a SMBH, an episode of rapid star formation, i.e. a starburst, is triggered in the vicinity of the SMBH (e.g. Silk, 2005; Levin, 2007; Carilli and Walter, 2013; Heckman and Best, 2014; Mercedes-Feliz et al., 2023). Moreover, the reservoir of baryons in the vicinity of the SMDS is replenished now via the merger. Thus, the SMBH can continue to grow at the Eddington rate, and be observed as the X-ray portion of the UHZ1 radiation. The stellar population embedded in those two merging halos is responsible for the majority of the IR observed data. In Fig. 7 we present a schematic representation of the scenario discussed above. Or, there could already be one halo in which the SMDS has already collapsed to a SMBH, and that halo then merges with another. The merger then creates a starburst in addition to the SMBH that was already there.